# appendices

### v2

@@ Abstract
Access to detailed war impact assessments is crucial for humanitarian organizations to effectively assist populations most affected by armed conflicts. However, maintaining a comprehensive understanding of the situation on the ground is challenging, especially in conflicts that cover vast territories and extend over long periods. This study presents a scalable and transferable method for estimating war-induced damage to buildings. We first train a machine learning model to output pixel-wise probability of destruction from Synthetic Aperture Radar (SAR) satellite image time series, leveraging existing, manual damage assessments as ground truth and cloud-based geospatial analysis tools for large-scale inference. We further post-process these assessments using open building footprints to obtain a final damage estimate per building. We introduce an accessible, open-source tool that allows users to adjust the confidence interval based on their specific requirements and use cases. Our approach enables humanitarian organizations and other actors to rapidly screen large geographic regions for war impacts. We provide two publicly accessible dashboards: a Ukraine Damage Explorer to dynamically view our pre-computed estimates, and a Rapid Damage Mapping Tool to easily run our method and produce custom maps.
Figure 1: Percentage of buildings likely damaged within the first two years of the war, aggregated by hromadas. The predictions were thresholded at 0.655, and only buildings larger than 50 m2 were considered.
The Russian invasion of Ukraine in February 2022 escalated a simmering conflict, which had previously seen unofficial Russian involvement limited to the country’s eastern regions, into a full-scale war between the two countries. Russian troops crossed into Ukraine from multiple fronts and advanced through the country, reorienting their geographical focus as Ukrainian forces mounted an unexpectedly fierce resistance. As the Russian troops advanced, they subjected numerous cities to heavy shelling and destroyed critical infrastructure. Millions of Ukrainians fled their homes as damage and destruction spread throughout the country. Two years into the conflict, the war has caused hundreds of thousands of casualties and inflicted billions of dollars in damages to Ukraine’s infrastructure [1, 2].
@@
Nowadays, the analysis of satellite imagery with machine learning offers opportunities to automate parts of the labor-intensive remote monitoring process. Multiple studies have demonstrated the efficiency and effectiveness of this approach for assessing building damage. The most accurate methods currently involve applying deep neural networks to VHR optical data [3, 4, 5, 6, 7, 8], often leveraging large dedicated datasets [9]. These approaches primarily focus on natural disasters, which typically require a one-time analysis of relatively localized, clustered damages. In contrast, armed conflicts can last for months, years, or even decades, resulting in spatially dispersed and gradually increasing destruction. Consequently, they require continuous monitoring over extended periods, so that automated change detection can on the one hand bring larger benefits, but on the other hand, is more challenging  [10].
@@
Recent work has begun to apply research specifically to the task of assessing building destruction induced by armed conflicts. In a seminal study, Mueller et al. [11] trained a CNN with VHR optical imagery and labels from UNOSAT [12] to assess war-induced building destruction in Syria, leveraging the persistent nature of war damage over time due to the absence of reconstruction efforts amidst ongoing fighting. However, accessing VHR optical imagery for conflict zones is costly, as commercial sources typically do not provide free access to their database in these situations [13], unlike after natural disasters [14, 15]. Despite being an active field of research [16, 17], relying on repeated, large-area coverage with VHR images to monitor armed conflicts at scale is not realistic for most actors. Very recently, Hou et al. [18] proposed a deep learning-based approach that can handle both VHR and moderate-resolution optical imagery.
@@
Beyond optical images, Synthetic Aperture Radar (SAR) imagery offers a promising alternative. SAR is an active sensor system that illuminates the Earth’s surface with microwave pulses and captures the backscattered signals. Unlike its optical counterpart, SAR can operate at any time of day (night, respectively) and is largely unaffected by clouds. SAR has a well-established history in damage assessment for both natural and anthropogenic events [19, 20]. Broadly, SAR-based change detection can be divided into incoherent and coherent methods. Incoherent approaches rely only on backscatter amplitude, typically comparing intensity or correlation differences between pre- and post-event images [21, 22, 23, 24, 25]. Coherent methods, on the other hand, use both amplitude and phase information through interferometry (InSAR), generally resulting in more accurate results [26, 27, 28]. Additionally, SAR’s sub-meter wavelength allows coherent techniques to detect changes far smaller than the pixel resolution, particularly when using permanent scatterers, which maintain coherence over long time series [29]. Coherent change detection is a standard approach in the InSAR community to assess building damage after a disaster rapidly [30, 31, 32, 33], c.f. the popular (and patented) Damage Proxy Maps. Recent works have combined machine learning and coherence data to produce more accurate maps, either with handcrafted features [34, 35] or by using recurrent neural networks trained on coherence time series [36, 37]. On the downside, generating high-quality coherence maps is complex and involves a loss of resolution. In this work, we focus on large-scale deployment on cloud-based platforms such as Google Earth Engine (GEE) [38], which do not support the Sentinel-1 SLC product [39] necessary for phase-based methods. We therefore opted for a purely amplitude-based approach.
@@
A few studies have already explored satellite-based damage assessment for the Ukraine War, but most focused on specific areas or events. Aimaiti et al. achieved 58% accuracy on UNOSAT labels in the Kyiv region based on handcrafted features from Sentinel-1 and Sentinel-2 [40]. Huang et al. used optical and coherence SAR time series to identify burnt areas, respectively urban destruction in Mariupol, reaching an overall accuracy of $\approx$60across four classes relative to labels retrieved from social media images [41]. Tavakkoliestahbanati et al. leverage multi-temporal InSAR to monitor deformation from a single event, the Kakhovka dam collapse [42]. Ballinger’s method is arguably the one most directly comparable to ours. It relies on a pixel-wise statistical $t$-test on Sentinel-1 time series, implemented directly in GEE. UNOSAT labels from various locations were then used to calibrate the decision threshold [43]. Lastly, Scher and Van Den Hoek have generated nationwide coherence maps and performed coherent change detection against a reference pre-invasion image. At the time of writing their maps have not yet been made public [44].
Open-Source Tool for Mapping War Impact
@@
In this work, we introduce an easy-to-use, open-source war impact mapping tool based purely on SAR amplitude data and demonstrate its effectiveness through a comprehensive building damage assessment across Ukraine. Our solution leverages existing, point-wise damage maps from UNOSAT [12] and public, open SAR imagery from the Sentinel-1 mission [39]. Harnessing the parallel computing capabilities of GEE [38], we train a model to estimate the likelihood of war-related changes from paired time series of Sentinel-1 backscatter. We choose a fixed one-year interval in 2020 as our reference period and iteratively generate damage likelihood maps for 3-month periods ranging from February 2021 up to February 2024. Our model, based on the Random Forest algorithm, allows for straightforward deployment on GEE and outputs damage maps with 10$\times$10 m grid spacing, the original ground sampling distance (GSD) of the Sentinel-1 GRD product. In a second processing step, we intersect the per-pixel maps with publicly available building footprints from Overture Maps [45] to generate a damage estimate per building. Our tool produces maps of war damage at a very low cost and enables continuous country-scale monitoring. These maps can serve both as a basis for large-scale, aggregated assessments and as guidance to focus manual verification efforts for individual settlements or buildings. We offer two public online dashboards: In the first one, users can dynamically view and explore the pre-computed damage maps described in the following. The second one allows users to run our method themselves for their desired locations and time periods and create custom maps.
@@
The key advantages of our approach are its scalability and ability to transfer to different geographic contexts. Using commercial VHR imagery for regular screening across entire conflict zones is not viable for humanitarian organizations [46], because it would be too resource-intensive, in terms of both direct data cost as well as infrastructure for data download, storage, and compute. By utilizing moderate-resolution satellite images that are free of charge and acquired with short, regular revisits, we enable a more sustainable monitoring approach. The use of SAR imagery has the additional advantage that the same method can be expected to work well in other conflict contexts, including those with different architectural patterns and/or with even more frequent cloud cover, as it is purely geometrical and does not rely on any location-specific a-priori assumptions. The proposed method is also fast, reproducible, and easy to adapt, even for users who are not experts in remote sensing or image analysis. We believe that our work could be a valuable and accessible resource and could be used to screen large conflict areas beyond our specific application case.
@@ Results
Figure 2: Building damage estimate for Chernihiv (standard confidence threshold 0.655), aggregated over the first two years of conflict. For clarity, we only present the damage heatmap in the main figure, while three zoomed insets show building-level predictions. Red dots represent UNOSAT annotations indicating buildings marked as either destroyed or severely damaged. The VHR satellite layer is displayed solely for visualization, all results are derived from 10 m-resolution Sentinel-1 images. Readers are encouraged to explore the maps on the interactive dashboard. Sources: Google Earth/Maxar satellite imagery, Overture Maps building footprints, UNITAR/UNOSAT damage annotations.
@@
Table 1: Quantitative detection performance relative to UNOSAT labels (at previously unseen locations), using a confidence threshold of 0.5. Values in parentheses indicate performance at the threshold of 0.655, reaching a target precision of 90%. The undamaged class is evaluated using the same locations, but acquisition windows before the full-scale invasion. All metrics are given as percentages.
We first evaluate our model’s predictions against additional damage assessments excluded from the training set (see Fig. 4 for the geographic distribution of our test set). The UNOSAT-labeled locations are compared to the corresponding raw pixel predictions of the model. True positive and false negative rates are counted in the 3-month periods from 2022 (after the invasion), and true negative and false positive rates are counted in the periods from 2021 (before the invasion). To guard against potential shifts between the point labels and the satellite tiles, e.g. due to high incidence angles or inaccurate annotation, we compare to the maximum predicted damage probability in a 3$\times$3 pixel window centered at the reference label. We report our results in Table 1. We obtain an F1-score of 74.9% for the damaged class and an AUC of 81.3%. Using the same evaluation protocol, we compare our result to the pixel-wise $t$-test (PWTT) method [43]. To our knowledge, this is the only method directly comparable to ours, since others either require coherence maps or assume that the precise date of the destruction event is known, which is not the case for our UNOSAT labels. Our method consistently outperforms that baseline, see Table 2. For details about the comparison see Appendix B.
@@
Table 2: Comparison of our method with the PWTT [43], with the latter calibrated for Ukraine. Metrics were computed on our test set, using the same settings for both approaches. All metrics are given in percentages.
The evaluation of held-out validation data also helps establish an appropriate threshold for distinguishing between damaged and undamaged pixels. Indeed, while the model has been trained on a balanced dataset, the actual distribution of damaged versus undamaged pixels is highly skewed, and the prevalence of undamaged pixels in the test set means that thresholding at damage probability 0.5 will cause a significant number of false alarms. To address this, we set a target precision of 90%, aligning with the reality of operational deployment, where it is crucial to keep the false positive rate at a manageable level. With this choice, we find an optimal threshold of 0.655. For details about the threshold optimization see Appendix C. With that value, the recall reaches 46.5%. We point out that there is an inherent trade-off between missed detections and false alarms, and the right threshold depends on the application task. Our dashboard allows the user to adjust it with a slider and explore how different thresholds affect the estimates. Results with the confidence threshold of 0.655 are also reported in Table 1.
Figure 3: Screenshot from the dashboard interface, displaying damage estimates for a region in Mariupol, alongside the Sentinel-1 time series visualizer.
@@
To illustrate the scalability of our approach, we have run damage assessments over the entire country. Our analysis indicates that over 400,000 buildings in Ukraine, or approximately 2.7% of all buildings, have likely sustained damage during the first two years of conflict. This estimate only accounts for buildings larger than 50 m2. Fig. 1 shows the percentage of buildings likely damaged within the first two years of the war, aggregated by hromadas, which is the finest administrative division in Ukraine that covers the entire territory. For the most impacted hromadas with over 10,000 buildings—Bakhmutska, Marinska, and Popasnianska—damage rates reach 42.6%, 29.0%, and 25.0%, respectively. We observe that this pattern of destruction correlates with the course of the war, the eastern side of the country having been a battlefield for months, while the western part has been relatively spared. For a subset of $\approx$1.8M buildings in our dataset, we were able to retrieve meta-data about the building function (such as ”residential” or ”medical”) from the OpenStreetMap catalog. Table 3 summarizes the damaged fractions for different building functions. In this preliminary analysis, we did not find any country-wide patterns that would suggest that certain building types are more prone to war damage than others. Note a more detailed study is needed to confirm or reverse this finding, as biases in the definition and distribution of the available OpenStreetMap building function labels could influence the result.
@@ Discussion
Dependence on External Data: Our model has specifically been trained on building data and therefore presupposes that settlement or building boundaries are available to mask out areas without human settlements (e.g., agriculture, forest, etc.), which would otherwise cause many false positives. We found that it is sufficient to utilize standard, global land cover maps [47], settlement layers [48], or building footprints. Still, the reliance on such external sources is a potential source of error (see Appendix E for possible issues specific to building footprints).
@@
Using Sentinel-1 data from GEE: Our study focuses on large-scale assessment and takes full advantage of GEE’s computational capabilities. However, it is important to note that while the Sentinel-1 data processed by GEE are orthorectified, they are not radiometrically terrain flattened. This limitation is particularly relevant in areas with significant slope. While most cities in Ukraine are on fairly flat terrain, our method may not perform as well in towns and cities located in more mountainous areas. Additionally, as previously mentioned, GEE can at present not ingest complex data, which prevents the use of phase coherence for change detection. In future research, we plan to explore alternative approaches, so as to fully exploit the potential of SAR data.
Reference Data: Our supervised learning approach necessitates a sufficient volume of reference data. Despite existing calls for large-scale datasets and open data programs dedicated to armed conflicts [13], such resources remain limited. Data availability varies significantly between different contexts: for some conflicts, e.g. the wars in Ukraine and Gaza, quite a lot of damage data is available; whereas there may be very little information about building damages in conflicts that receive less public attention. It is precisely in these under-reported conflicts that automated analysis could have the biggest impact. By offering an accessible open-access tool based on public satellite data, we hope to take one step toward a global monitoring solution. But to get there it will likely be necessary to collect at least some training data that represent the building patterns and imaging conditions of regions with under-reported conflicts.
Processing Time and User Experience: In principle, the proposed method is fast. Still, when running the web tool it may take several minutes to display the result, due to the technical overheads of GEE. As an alternative, one may download the corresponding image tiles and process them locally with the provided code. In the future, we hope to further accelerate the online tool.
@@
Post-processing the model outputs with additional data sources, as demonstrated for Ukraine with building footprints, adds an important dimension of information. Future research could seek to assimilate further sources of information, for instance, news reports or social media. Despite the inherent biases of such sources [49], they contain a wealth of complementary information that could be harvested, adding contextual knowledge and narrative depth to maps derived purely from satellite images. For example, attacks on residential buildings carry different policy implications if these buildings are inhabited than if civilians have fled an area in anticipation of strikes. Of course, even when data sources are combined to enhance the information derived from satellite images, contextual knowledge offered by observers and humanitarian organizations remains crucial for decision-making.

### v3

@@ Abstract
Access to detailed war impact assessments is crucial for humanitarian organizations to assist affected populations effectively. However, maintaining a comprehensive understanding of the situation on the ground is challenging, especially in widespread and prolonged conflicts. Here we present a scalable method for estimating building damage resulting from armed conflicts. By training a machine learning model on Synthetic Aperture Radar image time series, we generate probabilistic damage estimates at the building level, leveraging existing damage assessments and open building footprints. To allow large-scale inference and ensure accessibility, we tie our method to run on Google Earth Engine. Users can adjust confidence intervals to suit their needs, enabling rapid and flexible assessments of war-related damage across large areas. We provide two publicly accessible dashboards: a Ukraine Damage Explorer to dynamically view our precomputed estimates, and a Rapid Damage Mapping Tool to run our method and generate custom maps.
Figure 1: Spatial distribution of building damage in Ukraine after two years of conflict. Percentage of buildings likely damaged within the first two years of the war, aggregated by hromadas. The predictions were thresholded at 0.655, and only buildings larger than 50 m2 were considered.
Introduction
The Russian invasion of Ukraine in February 2022 escalated a simmering conflict, which had previously seen unofficial Russian involvement limited to the country’s eastern regions, into a full-scale war between the two countries. Russian troops crossed into Ukraine from multiple fronts and advanced through the country, reorienting their geographical focus as Ukrainian forces mounted an unexpectedly fierce resistance. As the Russian troops advanced, they subjected numerous cities to heavy shelling and destroyed critical infrastructure. Millions of Ukrainians fled their homes as damage and destruction spread throughout the country. Two years into the conflict, the war has caused hundreds of thousands of casualties and inflicted billions of dollars in damages to Ukraine’s infrastructure [kse_russia_will_pay, WorldBank2023].
@@ Introduction
Nowadays, the analysis of satellite imagery with machine learning offers opportunities to automate parts of the labor-intensive remote monitoring process. Multiple studies have demonstrated the efficiency and effectiveness of this approach for assessing building damage. The most accurate methods currently involve applying deep neural networks to VHR optical data [Gueguen2015, Xu2019, Lee2020, Zheng2021, Gholami2022, Kaur2023], often leveraging large dedicated datasets [xBD]. These approaches primarily focus on natural disasters, which typically require a one-time analysis of relatively localized, clustered damages. In contrast, armed conflicts can last for months, years, or even decades, resulting in spatially dispersed and gradually increasing destruction. Consequently, they require continuous monitoring over extended periods, so that automated change detection can on the one hand bring larger benefits, but on the other hand, is more challenging  [Sticher2024].
@@
Recent work has begun to apply research specifically to the task of assessing building destruction induced by armed conflicts. In a seminal study, Mueller2021 trained a CNN with VHR optical imagery and labels from UNOSAT [unosat] to assess war-induced building destruction in Syria, leveraging the persistent nature of war damage over time due to the absence of reconstruction efforts amidst ongoing fighting [Mueller2021]. However, accessing VHR optical imagery for conflict zones is costly, as commercial sources typically do not provide free access to their database in these situations [Bennett2022], unlike after natural disasters [MAXAROpen, PlanetDisaster]. Despite being an active field of research [Nabiee2022, Singh2023], relying on repeated, large-area coverage with VHR images to monitor armed conflicts at scale is not realistic for most actors. Very recently, Hou2024 proposed a deep learning-based approach that can handle both VHR and moderate-resolution optical imagery [Hou2024].
@@
Beyond optical images, Synthetic Aperture Radar (SAR) imagery offers a promising alternative. SAR is an active sensor system that illuminates the Earth’s surface with microwave pulses and captures the backscattered signals. Unlike its optical counterpart, SAR can operate at any time of day (night, respectively) and is largely unaffected by clouds. SAR has a well-established history in damage assessment for both natural and anthropogenic events [Plank2014, Ge2020]. Broadly, SAR-based change detection can be divided into incoherent and coherent methods. Incoherent approaches rely only on backscatter amplitude, typically comparing intensity or correlation differences between pre- and post-event images [Aoki1998, Matsuoka2004, Matsuoka2005, Liu2013, Uprety2013]. Coherent methods, on the other hand, use both amplitude and phase information through interferometry (InSAR), generally resulting in more accurate results [Matsuoka2000, Fielding2005, Gamba2007]. Additionally, SAR’s sub-meter wavelength allows coherent techniques to detect changes far smaller than the pixel resolution, particularly when using permanent scatterers, which maintain coherence over long time series [Ferretti2001]. Coherent change detection is a standard approach in the InSAR community to assess building damage after a disaster rapidly [Yun2015, Watanabe2016, Washaya2018, Tay2020], c.f. the popular (and patented) Damage Proxy Maps. Recent works have combined machine learning and coherence data to produce more accurate maps, either with handcrafted features [Mastro2022, Akhmadiya2022] or by using recurrent neural networks trained on coherence time series [Stephenson2022, Yang2024]. On the downside, generating high-quality coherence maps is complex and involves a loss of resolution. In this work, we focus on large-scale deployment and accessibility and opted for developing our approach on the cloud-based platform Google Earth Engine (GEE) [Gorelick2017]. Since GEE does not support the Sentinel-1 SLC product [sentinel_1] necessary for phase-based methods, our approach is purely amplitude-based.
@@
A few studies have already explored satellite-based damage assessment for the Ukraine War, but most focused on specific areas or events. Aimaiti2022 achieved 58% accuracy on UNOSAT labels in the Kyiv region based on handcrafted features from Sentinel-1 and Sentinel-2 [Aimaiti2022]. Huang2023 used optical and coherence SAR time series to identify burnt areas, respectively urban destruction in Mariupol, reaching an overall accuracy of $\approx$60across four classes relative to labels retrieved from social media images [Huang2023]. Tavakkoliestahbanati2024 leverage multi-temporal InSAR to monitor deformation from a single event, the Kakhovka dam collapse [Tavakkoliestahbanati2024]. Ballinger2024’s method is arguably the one most directly comparable to ours. It relies on a pixel-wise statistical $t$-test on Sentinel-1 time series, implemented directly in GEE. UNOSAT labels from various locations were then used to calibrate the decision threshold [Ballinger2024]. Lastly, Scher2023 have generated nationwide coherence maps and performed coherent change detection against a reference pre-invasion image. At the time of writing their maps have not yet been made public [Scher2023].
@@
In this work, we introduce an easy-to-use, open-source war impact mapping tool based purely on SAR amplitude data and demonstrate its effectiveness through a comprehensive building damage assessment across Ukraine. Our solution leverages existing, point-wise damage maps from UNOSAT [unosat] and public, open SAR imagery from the Sentinel-1 mission [sentinel_1]. Harnessing the parallel computing capabilities of GEE [Gorelick2017], we train a model to estimate the likelihood of war-related changes from paired time series of Sentinel-1 backscatter. We choose a fixed one-year interval in 2020 as our reference period and iteratively generate damage likelihood maps for 3-month periods ranging from February 2021 up to February 2024. Our model, based on the Random Forest algorithm, allows for straightforward deployment on GEE and outputs damage maps with 10$\times$10 m grid spacing, the original ground sampling distance (GSD) of the Sentinel-1 GRD product. In a second processing step, we intersect the per-pixel maps with publicly available building footprints from Overture Maps [OvertureMaps] to generate a damage estimate per building. Our tool produces maps of war damage at a very low cost and enables continuous country-scale monitoring. These maps can serve both as a basis for large-scale, aggregated assessments and as guidance to focus manual verification efforts for individual settlements or buildings. We offer two public online dashboards: In the first one, users can dynamically view and explore the pre-computed damage maps described in the following. The second one allows users to run our method themselves for their desired locations and time periods and create custom maps.
@@
The key advantages of our approach are its scalability, accessibility, and ability to transfer to different geographic contexts. Using commercial VHR imagery for regular screening across entire conflict zones is not viable for humanitarian organizations [Sticher2023], because it would be too resource-intensive, in terms of both direct data cost as well as infrastructure for data download, storage, and compute. By utilizing moderate-resolution satellite images that are free of charge and acquired with short, regular revisits, we enable a more sustainable monitoring approach. The use of SAR imagery has the additional advantage that the same method can be expected to work well in other conflict contexts, including those with different architectural patterns and/or with even more frequent cloud cover, as it is purely geometrical and does not rely on any location-specific a-priori assumptions. The proposed method is also fast, reproducible, and easy to adapt, even for users who are not experts in remote sensing or image analysis. We believe that our work could be a valuable and accessible resource and could be used to screen large conflict areas beyond our specific application case.
@@ Results
Figure 2: Damage estimates and UNOSAT labels for Chernihiv. Building damage estimate for Chernihiv thresholded with the standard confidence threshold of 0.655, aggregated over the first two years of conflict. For clarity, we only present the damage heatmap in the main figure, while three zoomed insets show building-level predictions. Red dots represent UNOSAT annotations indicating buildings marked as either destroyed or severely damaged. The VHR satellite layer is displayed solely for visualization, all results are derived from 10 m-resolution Sentinel-1 images. Readers are encouraged to explore the maps on the interactive dashboard. Sources: Google Earth/Maxar Technologies, Overture Maps building footprints, UNITAR/UNOSAT damage annotations.
@@
Table 1: Evaluation metrics. Quantitative detection performance relative to UNOSAT labels (at previously unseen locations), using a confidence threshold of 0.5. Values in parentheses indicate performance at the threshold of 0.655, reaching a target precision of 90%. The undamaged class is evaluated using the same locations, but acquisition windows before the full-scale invasion. All metrics are given as percentages.
We first evaluate our model’s predictions against additional damage assessments excluded from the training set (see Fig. 4 for the geographic distribution of our test set). The UNOSAT-labeled locations are compared to the corresponding raw pixel predictions of the model. True positive and false negative rates are counted in the 3-month periods from 2022 (after the invasion), and true negative and false positive rates are counted in the periods from 2021 (before the invasion). To guard against potential shifts between the point labels and the satellite tiles, e.g., due to high incidence angles or inaccurate annotation, we compare to the maximum predicted damage probability in a 3$\times$3 pixel window centered at the reference label. We report our results in Table 1. We obtain an F1-score of 74.9% for the damaged class and an AUC of 81.3%. Using the same evaluation protocol, we compare our result to the pixel-wise $t$-test (PWTT) method [Ballinger2024]. To our knowledge, this is the only method directly comparable to ours, since others either require coherence maps or assume that the precise date of the destruction event is known, which is not the case for our UNOSAT labels. Our method consistently outperforms that baseline, see Table 2. For details about the comparison see Supplementary Note 2.
@@
Table 2: Comparison with PWTT. Comparison of our method with the PWTT [Ballinger2024], with the latter calibrated for Ukraine. Metrics were computed on our test set, using the same settings for both approaches. All metrics are given in percentages.
The evaluation of held-out validation data also helps establish an appropriate threshold for distinguishing between damaged and undamaged pixels. Indeed, while the model has been trained on a balanced dataset, the actual distribution of damaged versus undamaged pixels is highly skewed, and the prevalence of undamaged pixels in the test set means that thresholding at damage probability 0.5 will cause a considerable number of false alarms. To address this, we set a target precision of 90%, aligning with the reality of operational deployment, where it is crucial to keep the false positive rate at a manageable level. With this choice, we find an optimal threshold of 0.655. For details about the threshold optimization see Supplementary Note 3. With that value, the recall reaches 46.5%. We point out that there is an inherent trade-off between missed detections and false alarms, and the right threshold depends on the application task. Our dashboard allows the user to adjust it with a slider and explore how different thresholds affect the estimates. Results with the confidence threshold of 0.655 are also reported in Table 1.
Figure 3: The Rapid Damage Mapping Tool. Screenshot from the dashboard interface, displaying damage estimates for a region in Mariupol, alongside the Sentinel-1 time series visualizer.
@@
To illustrate the scalability of our approach, we have run damage assessments over the entire country. Our analysis indicates that over 400,000 buildings in Ukraine, or approximately 2.7% of all buildings, have likely sustained damage during the first two years of conflict. This estimate only accounts for buildings larger than 50 m2. Fig. 1 shows the percentage of buildings likely damaged within the first two years of the war, aggregated by hromadas, the finest administrative division in Ukraine covering the entire territory. For an analysis of how damage has evolved over time, see Supplementary Note 4. For the most impacted hromadas with over 10,000 buildings—Bakhmutska, Marinska, and Popasnianska—damage rates reach 42.6%, 29.0%, and 25.0%, respectively. We observe that this pattern of destruction correlates with the course of the war, the eastern side of the country having been a battlefield for months, while the western part has been relatively spared. For a subset of $\approx$1.8M buildings in our dataset, we were able to retrieve meta-data about the building function (such as ”residential” or ”medical”) from the OpenStreetMap catalog. Table 3 summarizes the damaged fractions for different building functions. In this preliminary analysis, we did not find any country-wide patterns that would suggest that certain building types are more prone to war damage than others. Note a more detailed study is needed to confirm or reverse this finding, as biases in the definition and distribution of the available OpenStreetMap building function labels could influence the result.
@@ Discussion
Google Earth Engine: Our study focuses on large-scale assessment and accessibility, leveraging GEE’s computational capabilities. However, this reliance also imposes significant constraints. First, it restricts the choice of machine learning methods, as only a limited set of classification algorithms are available at the time of writing. Second, while Sentinel-1 data processed by GEE is orthorectified, it is not radiometrically terrain-flattened, which can introduce distortions in areas with pronounced slopes. Although most cities in Ukraine are situated on relatively flat terrain, our method may be less reliable in towns and cities located in mountainous regions. Additionally, as previously mentioned, GEE currently does not support the ingestion of complex SAR data, preventing the use of phase coherence for change detection. Lastly, while the proposed method is computationally efficient in principle, the web tool may take several minutes to generate results due to GEE’s technical overhead. As an alternative, users can download the relevant image tiles and process them locally using the provided code. In the future, we hope to further accelerate the online tool.
Dependence on External Data: Our model has specifically been trained on building data and therefore presupposes that settlement or building boundaries are available to mask out areas without human settlements (e.g., agriculture, forest, etc.), which would otherwise cause many false positives. We found that it is sufficient to utilize standard, global land cover maps [Brown2022], settlement layers [gshl], or building footprints. Still, the reliance on such external sources is a potential source of error (see Supplementary Note 5 for possible issues specific to building footprints).
@@
Reference Data: Our supervised learning approach necessitates a sufficient volume of reference data. Despite existing calls for large-scale datasets and open data programs dedicated to armed conflicts [Bennett2022], such resources remain limited. Data availability varies considerably between different contexts: for some conflicts, e.g. the wars in Ukraine and Gaza, quite a lot of damage data is available; whereas there may be very little information about building damages in conflicts that receive less public attention. It is precisely in these under-reported conflicts that automated analysis could have the biggest impact. By offering an accessible open-access tool based on public satellite data, we hope to take one step toward a global monitoring solution. But to get there it will likely be necessary to collect at least some training data that represent the building patterns and imaging conditions of regions with under-reported conflicts.
@@
Post-processing the model outputs with additional data sources, as demonstrated for Ukraine with building footprints, adds an important dimension of information. Future research could seek to assimilate further sources of information, for instance, news reports or social media. Despite the inherent biases of such sources [Miller2022], they contain a wealth of complementary information that could be harvested, adding contextual knowledge and narrative depth to maps derived purely from satellite images. For example, attacks on residential buildings carry different policy implications if these buildings are inhabited than if civilians have fled an area in anticipation of strikes. Of course, even when data sources are combined to enhance the information derived from satellite images, contextual knowledge offered by observers and humanitarian organizations remains crucial for decision-making.